### v2

# An analysis of nuclear parton distribution function based on Kullback-Leibler divergence

Abstract
In this work, we propose a method to quantify the difference between nuclear parton distribution functions (nPDFs) in different nuclei and parton distribution functions (PDFs) in free nucleons using the Kullback-Leibler (KL) divergence, a measure widely employed in quantum information theory. By introducing certain constraints and the “minimum relative entropy” hypothesis, we can determine the shape of the structure function in the intermediate-$x$ region, which is intimately connected with the renowned EMC effect. For quark nPDFs, our results align with the latest global fits to experimental data. This agreement suggests that the KL divergence-based methodology may provide novel insight into the structure of nucleons, particularly in cases where experimental data and theoretical QCD constraints are limited, such as those pertinent to gluon nPDFs. Therefore, we applied this methodology to gluon nPDFs, analyzing the results from two commonly used global fitting groups, EPPS21 and nNNPDF3.0. Our analysis reveals that the EPPS21 results align more closely with the “minimum relative entropy” hypothesis. This finding underscores the utility of the proposed method and provides a valuable reference for future global fitting of nPDFs.
I Introduction
Parton distribution functions (PDFs) are a kind of quantum correlation functions which characterize the probability that the parton carries a certain momentum fraction of the hadron momentum. The PDFs serve as fundamental inputs for precision Standard Model (SM) baseline predictions at the LHC and other collider facilities, thus their determination is a key focus in high-energy physics Hou:2019efy ; Bailey:2020ooq ; Lai:2010vv ; Martin:2009iq ; Alekhin:2012ig ; Barone:2000tx .
When free nucleons become bound within atomic nuclei, nuclear parton distribution functions (nPDFs) become necessary. One of the primary goals of nuclear physics is achieving a comprehensive description of the structure of atomic nuclei. The nPDFs, which is responsible for the motion of partons in the nuclear medium, is one of the most poorly understood parts of nuclear structure. A prime example is the unexpected EMC effect, which indicates the quark PDFs in bound nucleons are modified compared with free ones in the intermediate-$x$ region EuropeanMuon:1983wih ; NewMuonNMC:1990xyw ; Gomez:1993ri ; Seely:2009gt . The complex nature of nonperturbative QCD renders the analytical solution of nPDFs unattainable. Furthermore, the lattice QCD also remains impractical on this issue, as nuclear structure represents a quintessential many-body problem that currently exceeds the computational capabilities and resources available for lattice simulations. Even for the very light nucleus, current lattice calculations remain at an exploratory stage Detmold:2020snb ; Amarasinghe:2021lqa ; Chen:2024rgi . Therefore, nearly all current available nPDFs are extracted from global fitting analysis of experimental data Eskola:2021nhw ; AbdulKhalek:2022fyi ; Duwentaster:2022kpv ; Helenius:2021tof .
Moreover, our knowledge of gluon nPDFs is considerably more limited. While modern global fitting analyses yield increasingly precise quark nPDFs—with nuclear modifications for quarks in the intermediate-to-large $x$ region being broadly consistent across different fitting groups—the gluon nPDF remains substantially more uncertain. In contrast to the quark sector, predictions for gluons from different groups often differ markedly. Therefore, developing new methodologies to pin down parton (especially gluon) nPDFs is a significant scientific goal.
The recent rapid development of quantum information theory has drawn attention in high-energy physics for its potential applications to long-standing problems, possibly revealing fundamental connections to underlying physical mechanism Galindo:2001ei ; ELEMENTSOFINFORMATIONTHEORY . Quantum information theory has introduced novel observables—various types of entropies that enable an extended characterization of hadron structure or scattering processes beyond classical approaches. Studies in this area have proliferated significantly Wang:2014lua ; Kharzeev:2017qzs ; Beane:2018oxh ; Gotsman:2020bjc ; Fedida:2022izl ; Hentschinski:2022rsa ; Liu:2022grf ; Han:2020vjp ; Ma:2018wtw ; Hagiwara:2017uaz ; Vedral:2002zz ; Nam:2025hei ; Miller:2023ujx ; Zhang:2023oja ; Bloss:2025ywh ; Hu:2025jne ; Wang:2025inr ; Lin:2025eci . Here, we want to find an appropriate entropy for the investigation of nPDFs. First, both PDF and nPDF are probability distributions. Second, we want to quantify the difference between these two distributions. The appropriate quantity is the KL divergence $\int p(x)\ln(p(x)/q(x))dx$ (also known as relative entropy) Kullback:1951zyt ; Wehrl:1978zz ; Benito-Calvino:2022kqa ; Cao:2022iqh ; Cao:2022ajt , which is a type of statistical distance: a measure of how much a reference distribution $q(x)$ is different from a true distribution $p(x)$. If we treat the PDF as the reference distribution and the nPDF as the true distribution, their ratio reflects the nuclear modification, whose fundamental mechanism may be recast in the quantum-theoretic terms of KL divergence. When a free nucleon is placed inside an atomic nucleus, its PDFs are modified into nPDFs due to interactions with surrounding nucleons. Understanding the precise mechanism of this transformation remains a challenge. This paper explores an information-theoretic approach to this problem, elaborating on how concepts such as the “minimum relative entropy” hypothesis can shed new light on this transformation.
In this paper, we quantify the difference between PDFs and nPDFs by KL divergence and determine the quark nPDFs in the intermediate-$x$ region (EMC region) through minimum relative entropy hypothesis, which proposes that when a free nucleon is bound in a nucleus, its PDFs transform into nPDFs in a way that minimizes the relative entropy between them, subject to known physical constraints. This approach frames the transformation as an optimization problem with fixed boundary conditions, providing a principle to determine shape of nPDFs. In turn, this methodology offers a new benchmark for assessing the quality of global fits to gluon nPDFs, which will be elaborated below.
The organization of this paper is as follows. In Sec. II, we give the formalism for structure functions in deeply inelastic scattering (DIS). Sec. III presents the definition of KL divergence and discusses its properties. In Sec. IV, we develop a framework and present results for quantifying and determining quark structure function using the KL divergence. We turn to the investigation of gluon nPDFs in Sec. V. We finish with the conclusion in Sec. VI.
II Formalism for structure functions in DIS
The cross sections for neutral-current ($eN\to eX$) DIS on unpolarized nucleons can be written in terms of the structure functions
$\frac{d^{2}\sigma_{\textrm{DIS}}}{dxdy}\!=\!\frac{4\pi\alpha^{2}}{Q^{4}}s\Big((1-y)F_{2}(x,Q^{2})+y^{2}xF_{1}(x,Q^{2})\Big)\,$, (1)
here $x$ is the Bjorken scaling variable, $y$ denotes the inelasticity, $Q^{2}$ represents the negative square of four-momentum transfer from lepton to hadron, and $s$ stands for the square of the center-of-mass energy. It should be noted that Eq. (1) only considers the neutral-current photon exchange at leading order (LO) in the strong coupling constant $\alpha_{s}$, and this equation can be further reduced by Callan-Gross relation $F_{2}=2xF_{1}$ Callan:1969uq
$\frac{d^{2}\sigma_{\textrm{DIS}}}{dxdy}=\frac{2\pi\alpha^{2}}{Q^{4}}s\Big(1+(1-y)^{2}\Big)F_{2}(x,Q^{2})\,.$ (2)
This simplified form in Eq. (2) is adopted in this study to facilitate the subsequent analysis using the KL divergence. At LO, the connection between the structure function $F_{2}$ and the quark distributions is given by
$F_{2}(x,Q^{2})=\sum_{q}e_{q}^{2}\,x\Big[f_{q}(x,Q^{2})+f_{\bar{q}}(x,Q^{2})\Big]\,$, (3)
where $e_{q}$ denotes the quark electric charge and $f_{q}(\bar{f}_{q})$ represents the PDF of (anti)parton $q$. When considering the case of atomic nucleus, the structure function becomes
$F_{2}^{A}(x,Q^{2})/A$ $=$ $\sum_{q}e_{q}^{2}\,x\Big[\frac{Z}{A}\left(f_{q}^{p/A}(x,Q^{2})+f_{\bar{q}}^{p/A}(x,Q^{2})\right)$ (4)
$+\frac{N}{A}\left(f_{q}^{n/A}(x,Q^{2})+f_{\bar{q}}^{n/A}(x,Q^{2})\right)\Big]\,$,
where $Z$ is the atomic number, $N$ is the neutron number, and $A$ is the mass number of the nucleus. $f_{q}^{p/A}$ represents the nPDFs of parton $q$ in a proton bound in nucleus A, and $f_{q}^{n/A}$ corresponds to a neutron.
We adopted the following notations for concision,
$\left\{\begin{aligned} &f_{q}^{A}(x,Q^{2})=\frac{Z}{A}f_{q}^{p/A}(x,Q^{2})+\frac{N}{A}f_{q}^{n/A}(x,Q^{2})\,,\\ \\ &f_{\bar{q}}^{A}(x,Q^{2})=\frac{Z}{A}f_{\bar{q}}^{p/A}(x,Q^{2})+\frac{N}{A}f_{\bar{q}}^{n/A}(x,Q^{2})\,.\end{aligned}\right$. (5)
Therefore, the structure function for nucleus A can be simplified
$F_{2}^{A}(x,Q^{2})/A$ $=$ $\sum_{q}e_{q}^{2}\,x\Big[f_{q}^{A}(x,Q^{2})+f_{\bar{q}}^{A}(x,Q^{2})\Big]\,.$ (6)
Numerous experimental and theoretical studies have been dedicated to describing this structure function ParticleDataGroup:2024cfk ; collinsbook ; Hen:2016kwk ; Radyushkin:2022qvt ; Zhang:2018diq ; Frankfurt:2011cs ; Chen:2020ody ; Chen:2015tca ; Wei:2016far . Although we do know lots of information about it, we still cannot obtain the exact distributions from first principle. The currently available and best results come from global fits of experimental data. For instance, the EPPS21 Collaboration presents the numerical values of $f_{q}^{A}$ and $f_{\bar{q}}^{A}$ datasetofEPPS21 , which show good agreement with experimental results NewMuon:1995cua ; NewMuon:1995tgs ; NewMuon:1996yuf ; NewMuon:1996gam ; EuropeanMuon:1992pyr .
The EMC effect describes a deviation from unity in the ratio of per-nucleon deep inelastic structure functions for nucleus A to deuteron $(F_{2}^{A}(x,Q^{2})/A)/(F_{2}^{d}(x,Q^{2})/2)$ in the intermediate-$x$ region, which in turn may leave a signature on quantum information-theoretic quantities such as the KL divergence, and serve as a proxy for the nuclear modifications to the nucleon structure.
III The Kullback-Leibler divergence for structure function
In this work, we deploy the KL divergence between PDFs and nPDFs to assess the nuclear modifications. For the sake of being self-contained, we start with a short introduction of the KL divergence. Mathematically, the KL divergence between two probability distributions is defined as
$D_{\mathrm{KL}}(p\|q)=\int p(x)\ln\frac{p(x)}{q(x)}dx\,.$ (7)
From here on, we will denote $q(x)$ as the reference distribution and $p(x)$ as the true distribution. Here we employ the KL divergence, which is also referred to as relative entropy in literature, with units of nats (with the base of the logarithm $e$).
KL divergence is a fundamental quantity in probability and quantum information theory, it arises as an expected logarithm of the likelihood ratio of two distributions $p(x)$ and $q(x)$. In the following, we outline its distinguishing features that the KL divergence is a measure of the “distance” between the reference probability function $q(x)$ and the true probability function $p(x)$. It is always a non-negative real number, with value 0 if and only if the reference function is identical to the true function. The nonnegativity of KL divergence implicitly assumes a prerequisite condition: both functions $q(x)$ and $p(x)$ are normalized to unity, $\int q(x)dx=1,\,\int p(x)dx=1$.
As discussed above, KL divergence is advantageous in analyzing nPDFs from PDFs, and it gives us a quantitative measure of the information needed to describe the true distribution (nPDFs) when starting from the reference scenario (PDFs). Having a feeling for the KL divergence, we turn to its physical application on quantifying and determining the nPDFs of quarks and gluons in the subsequent sections.
IV Framework and results for quark nPDFs
In this section, we provide calculation method of the KL divergence for different nuclei. The EMC experiment measured the ratio of per-nucleon structure functions for nucleus A to deuteron $(F_{2}^{A}(x,Q^{2})/A)/(F_{2}^{d}(x,Q^{2})/2)$. Theoretically, a major advantage of employing this ratio lies in its insensitivity to higher-order perturbative QCD corrections, as the perturbative contributions are largely irrelevant to the nonperturbative PDFs and nPDFs. Here, we take the numerator of this ratio as the true distribution $p^{A}(x)$ and the denominator as the reference distribution $q^{d}(x)$. In this work, $Q^{2}$ is set to $10\,\text{GeV}^{2}$, this is a typical scale in DIS processes. To keep the notation concise, the explicit $Q^{2}$ dependence is implicitly omitted.
For the reference distribution, deuteron is regarded as an approximation to a free proton and neutron system since it is loosely bound and the average distance between the nucleons is large. Therefore, we adopt the CT18A free PDFs to represent the nucleon nPDFs within the deuteron Hou:2019efy . For the true distribution, we take values from the EPPS21 parametrization for different kinds of nuclei  [^1] https://research.hip.fi/qcdtheory/nuclear-pdfs/epps21/ Eskola:2021nhw . Since we perform numerical integration of global fitting results, we must renounce to compute the KL divergence across the entire $x$-range $x\in[0,1]$. Here, the range for numerical integration is $x\in[10^{-6},0.99]$, a technical problem arises accordingly is the normalization. As previously introduced, the nonnegativity of KL divergence inherently requires that both distributions be individually normalized to unity. Otherwise, the calculated KL divergence could yield negative values, which is very counterintuitive for a measure representing any sort of entropy. Hence we introduce two normalization factors $z_{p}$ and $z_{q}$, the normalized distributions are
$\left\{\begin{aligned} &\left.p^{A}(x)\right|_{\textrm{nor}}=\frac{1}{\int_{10^{-6}}^{0.99}p^{A}(x)dx}\,p^{A}(x)=z_{p}\,p^{A}(x)\,,\\ &\left.q^{d}(x)\right|_{\textrm{nor}}=\frac{1}{\int_{10^{-6}}^{0.99}q^{d}(x)dx}\,q^{d}(x)=z_{q}\,q^{d}(x)\,.\end{aligned}\right$. (8)
The KL divergence between the structure functions of nucleus A and deuteron is calculated through
$D_{\mathrm{KL}}^{\textrm{full}}(p^{A}\|q^{d})=\int_{10^{-6}}^{0.99}\left.p^{A}(x)\right|_{\textrm{nor}}\ln\frac{\left.p^{A}(x)\right|_{\textrm{nor}}}{\left.q^{d}(x)\right|_{\textrm{nor}}}\,dx\,.$ (9)
Moreover, we are particularly interested in the nucleon structure function in the intermediate-$x$ region, as it is closely related to the EMC effect. Experimental data reveal that the ratio of per-nucleon structure functions for nucleus A to deuteron can be divided into four distinct regions: shadowing region, antishadowing region, EMC region, and Fermi motion region. In the EMC region, the ratio monotonically decreases from a maximum to a minimum. Within the EPPS21 set, this region spans $x\in[0.25,0.65]$. We have also computed the KL divergence within this region using the same method described above, except that the integration limits in Eqs. (8) and (9) have been changed from $10^{-6},0.99]$ to $[0.25,0.65]$. More details about these resulting KL divergences for different nuclei are collected in Appendix A.
These results essentially quantify the findings which have already been known. More importantly, we are interested in whether the KL divergence can provide new insights into nucleon structure, especially in the EMC region. Prior to the discovery of the EMC effect, it was commonly believed that the ratio $(F_{2}^{A}(x)/A)/(F_{2}^{d}(x)/2)$ would be unity in the intermediate-$x$ region, implying a corresponding KL divergence of zero. However, the EMC effect revealed that this ratio exhibits a deviation from unity, showing a monotonic decrease in the approximate $x$-range of 0.25 to 0.65.
For a given nucleus A, if we fix the two endpoints at $x=0.25$ and $x=0.65$, respectively, a question arises: can we establish a criterion to determine the shape of the structure function within the EMC region? This seems difficult since we are trying to construct this function with insufficient data—neither perturbative QCD nor lattice QCD can be applied on this issue. Interestingly, the problem encountered here bears a striking resemblance to the famous brachistochrone problem proposed by Bernoulli in 1696 brachistochronecurve . Both involve finding the extremum of a functional under fixed boundary conditions: the brachistochrone problem seeks the minimal time of descent, while our task requires establishing an extremum principle for a physical observable.
The physical observable we choose is the KL divergence utilized above,
$D_{\mathrm{KL}}^{\textrm{EMC}}(\bar{p}^{A}\|q^{d})=\int_{0.25}^{0.65}\left.\bar{p}^{A}(x)\right|_{\textrm{nor}}\ln\frac{\left.\bar{p}^{A}(x)\right|_{\textrm{nor}}}{\left.q^{d}(x)\right|_{\textrm{nor}}}\,dx\,.$ (10)
We assume that, with the endpoints fixed, the actual structure function of the nucleus tends to minimize the KL divergence in Eq. (10), which we refer to as the “minimum relative entropy” hypothesis. It is important to emphasize that this hypothesis serves a role analogous to the shortest time principle in the brachistochrone problem. With this hypothesis, the structure function can be computed and subsequently compared with the results from global fitting analyses. The $\bar{p}^{A}(x)$ in Eq. (10) represents the structure function we aim to determine. First, we adopt a polynomial form of parametrization
$\bar{p}^{A}_{\textrm{poly}}(x)=\big(a_{1}x^{3}+b_{1}x^{2}+c_{1}x+d_{1}\big)\,q^{d}(x)\,.$ (11)
Using the EPPS21 global fits, we can derive constraints at the endpoints, which determine two parameters in Eq. (11), while the remaining two are fixed by the minimum relative entropy hypothesis. In Figure 1, we provide two-dimensional visualization of the KL divergence as a function of $a_{1}$ and $b_{1}$ parameters. The KL divergence forms a distribution in the parameter space that allows us to pin down its minimum point. The position marked with a red dot indicates the corresponding $a_{1}$ and $b_{1}$ with minimum KL divergence. The black dot represents the result of EPPS21 global analyses. These two points are found to be in close proximity within the parameter space. To further illustrate the effect of deviating from the minimum relative entropy solution, two additional dots, shown in magenta and orange, are chosen to serve as representative examples. The structure functions corresponding to all these four dots will be presented in Figure 2.
Figure 1: The KL divergence as a function of $a_{1}$ and $b_{1}$. The red dot indicates the position with minimum KL divergence for 4He, 12C, 56Fe and 208Pb, at the momentum transfer $Q^{2}=10\,{\rm GeV}^{2}$. The orange and magenta dots mark two representative reference solutions, which will be used for comparisons in Figure 2.
Figure 2: The obtained structure function $\bar{p}^{A}_{\textrm{poly}}(x)$ with minimum KL divergence (red line), the structure function $p^{A}(x)$ given by the global fitting of experimental data (black line), and the two reference structure functions $\bar{p}^{A}_{\textrm{ref1}}(x)$ and $\bar{p}^{A}_{\textrm{ref2}}(x)$ (orange and magenta lines).
We present the explicit results of the polynomial parametrization for 4He, 12C, 56Fe and 208Pb,
$\bar{p}^{{}^{4}\textrm{He}}_{\textrm{poly}}(x)$ $\!=\!$ $\big(\!-\!5.89\,x^{3}\!+\!7.39\,x^{2}\!-\!3.00\,x\!+\!1.38\big)\,q^{d}(x),$
$\bar{p}^{{}^{12}\textrm{C}}_{\textrm{poly}}(x)$ $\!=\!$ $\big(\!-\!7.60\,x^{3}\!+\!9.53\,x^{2}\!-\!3.86\,x\!+\!1.48\big)\,q^{d}(x),$
$\bar{p}^{{}^{56}\textrm{Fe}}_{\textrm{poly}}(x)$ $\!=\!$ $\big(\!-\!11.51\,x^{3}\!+\!14.43\,x^{2}\!-\!5.85\,x\!+\!1.72\big)\,q^{d}(x),$
$\bar{p}^{{}^{208}\textrm{Pb}}_{\textrm{poly}}(x)$ $\!=\!$ $\big(\!-\!16.23\,x^{3}\!+\!20.32\,x^{2}\!-\!8.22\,x\!+\!1.99\big)\,q^{d}(x).$
Appendix B lists the four parameters ($a_{1}$, $b_{1}$, $c_{1}$ and $d_{1}$) in Eq. (11). Therein $c_{1}$ and $d_{1}$ are determined by the constraints at the two endpoints, while $a_{1}$ and $b_{1}$ are obtained by minimizing the KL divergences in Eq. (10) for different nuclei.
With the results in Eq. (IV), we can plot the structure function given by the minimum relative entropy hypothesis. Figure 2 shows a comparison between them and the global fitting results from EPPS21 set. Additionally, for reference, we also plot two curves which yield larger KL divergences, corresponding to the orange and magenta dots in Figure 1, respectively. From Figure 2, it can be observed that the structure functions given by Eq. (IV) are rather close to the experimental data. This indicates the minimum relative entropy hypothesis can yield reasonable nuclear structure functions (or the quark nPDFs), consistent with the popular nPDFs from QCD global analysis.
In addition, regarding the dependence on the functional form of the parametrization. A reasonable test is necessary. In Eq. (11), we adopted a polynomial form and expanded it to the cubic term $x^{3}$. We have also extended the expansion to the quartic term $x^{4}$, the resulting $\bar{p}^{A}_{\textrm{poly}}(x)$ remained identical to the cubic case. Therefore, we have not shown them for the sake of brevity. Moreover, the widely used canonical parametrization form with exponential and power-law endpoint behaviour were also tested,
$\bar{p}^{A}_{\textrm{can}}(x)=\big(a_{2}x^{b_{2}}(1-x)^{c_{2}}e^{d_{2}x}\big)\,q^{d}(x)\,.$ (13)
In line with the previous method, these parameters can be determined by the minimum relative entropy hypothesis,
$\bar{p}^{{}^{4}\textrm{He}}_{\textrm{can}}(x)$ $\!=\!$ $\big(0.20\,x^{-0.69}(1-x)^{1.39}e^{4.13x}\big)\,q^{d}(x),$
$\bar{p}^{{}^{12}\textrm{C}}_{\textrm{can}}(x)$ $\!=\!$ $\big(0.13\,x^{-0.90}(1-x)^{1.81}e^{5.37x}\big)\,q^{d}(x),$
$\bar{p}^{{}^{56}\textrm{Fe}}_{\textrm{can}}(x)$ $\!=\!$ $\big(0.039\,x^{-1.40}(1-x)^{2.84}e^{8.43x}\big)\,q^{d}(x),$
$\bar{p}^{{}^{208}\textrm{Pb}}_{\textrm{can}}(x)$ $\!=\!$ $\big(0.008\,x^{-2.09}(1-x)^{4.25}e^{12.57x}\big)\,q^{d}(x).$
Appendix B collects the values of the four parameters ($a_{2}$, $b_{2}$, $c_{2}$ and $d_{2}$) in Eq. (13).
We now summarize the key findings obtained so far. In Table 1, different KL divergences are compiled. The second column lists the KL divergence calculated using the results from the EPPS21 set, while the third and fourth column present the calculated minimum KL divergence from Eq. (IV) and Eq. (IV), respectively. As one can see, the obtained minimum KL divergences given by the polynomial and canonical parametrization forms show remarkable consistency with each other. This agreement becomes more visually discernible when displaying the structure functions together. Figure 3 shows these two kinds of structure functions obtained from the polynomial and canonical parameterizations, as well as the global fitting results from EPPS21 set. Visually indistinguishable differences are observed between the structure functions obtained from these two parameterizations.
These results demonstrate the robustness of the calculated minimum KL divergence across different parametrization forms. When the two endpoints in the intermediate-$x$ region are fixed, both the value of the minimum KL divergence and the shape of the nuclear structure function are almost exclusively determined by the minimum relative entropy hypothesis. It should be emphasized that, although we adopt two commonly used parametrizations of nPDFs, the values of their parameters are not determined by fitting. Instead, they are obtained by applying the minimum relative entropy hypothesis, which selects the parameters that minimize the relative entropy.
Table 1: The comparisons between the calculated KL divergence from EPPS21 global fitting nPDF sets and the calculated minimum KL divergences with the polynomial as well as canonical parameterizations in intermediate-$x$ region $[0.25,0.65]$ for different nuclei.
   Nucleus   $D_{\mathrm{KL}}^{\textrm{EMC}}(p^{A}\|q^{d})$   $D_{\mathrm{KL}}^{\textrm{EMC}}(\bar{p}^{A}_{\textrm{poly}}\|q^{d})$   $D_{\mathrm{KL}}^{\textrm{EMC}}(\bar{p}^{A}_{\textrm{can}}\|q^{d})$
3He $0.46\times 10^{-4}$ $1.3\times 10^{-5}$ $1.2\times 10^{-5}$
4He $2.32\times 10^{-4}$ $3.0\times 10^{-5}$ $2.9\times 10^{-5}$
9Be $4.75\times 10^{-4}$ $7.0\times 10^{-5}$ $6.9\times 10^{-5}$
12C $3.70\times 10^{-4}$ $5.0\times 10^{-5}$ $5.0\times 10^{-5}$
56Fe $8.17\times 10^{-4}$ $1.21\times 10^{-4}$ $1.21\times 10^{-4}$
197Au $16.19\times 10^{-4}$ $2.52\times 10^{-4}$ $2.58\times 10^{-4}$
208Pb $16.85\times 10^{-4}$ $2.63\times 10^{-4}$ $2.70\times 10^{-4}$
Figure 3: The obtained structure function $\bar{p}^{A}_{\textrm{poly}}(x)$ from polynomial form of parameterization (red line), the structure function $\bar{p}^{A}_{\textrm{can}}(x)$ from canonical form of parametrization (blue dashed line), as well as the structure function $p^{A}(x)$ given by the global fitting of experimental data (black line).
To quantify the perceived separation between various structure functions and the global fitting results given by EPPS21 in Figures 2 and 3, we make use of the $L^{2}$ norm which is defined as
$N(\bar{p}^{A}_{\textrm{poly}})=\sqrt{\int_{0.25}^{0.65}\Big(\bar{p}^{A}_{\textrm{poly}}(x)-p^{A}(x)\Big)^{2}dx}\,$, (15)
here $p^{A}(x)$ is the structure function from EPPS21. The results are listed in the second column of Table 2. It is evident that the norms calculated from $\bar{p}^{A}_{\textrm{poly}}(x)$ are much less than the reference curves $p^{A}_{\textrm{ref1}}(x)$ and $p^{A}_{\textrm{ref2}}(x)$, indicating the minimum relative entropy hypothesis could serve as a reliable criterion in constructing nPDFs of bound nucleon. Moreover, the two different parameterization forms (polynomial and canonical) yield nearly identical norms in Table 2.
These findings suggest that the discrepancy between the structure functions derived from the minimum relative entropy hypothesis and those obtained from global fits can be quantified by computing their $L^{2}$ norms. For the quark nPDFs considered in this section, the minimum relative entropy hypothesis produces values close to the global fits, resulting in a small $L^{2}$ norm (on the order of $10^{-3}$ to $10^{-4}$). In the following section, we adopt this norm-based comparison as a novel benchmark for evaluating the quality of global fits to gluon nPDFs.
Table 2: The norms of structure functions with minimum KL divergences (the second column) and norms of reference structure functions (the third and fourth columns).
   Nucleus   $N(\bar{p}^{A}_{\textrm{poly}})$ or $N(\bar{p}^{A}_{\textrm{can}})$    $N(\bar{p}^{A}_{\textrm{ref1}})$    $N(\bar{p}^{A}_{\textrm{ref2}})$
3He $6.61\times 10^{-4}$ $2.11\times 10^{-2}$ $2.33\times 10^{-2}$
4He $1.79\times 10^{-3}$ $50\times 10^{-2} 2.2$$25\times 10^{-2}$
9Be $2.38\times 10^{-3}$ $4.07\times 10^{-2}$ $2.91\times 10^{-2}$
12C $2.20\times 10^{-3}$ $49\times 10^{-2} 2.2$$26\times 10^{-2}$
56Fe $3.10\times 10^{-3}$ $1.22\times 10^{-2}$ $3.03\times 10^{-2}$
197Au $4.11\times 10^{-3}$ $1.34\times 10^{-2}$ $2.31\times 10^{-2}$
208Pb $4.14\times 10^{-3}$ $2.53\times 10^{-2}$ $2.89\times 10^{-2}$
V Determining gluon nPDFs through KL divergence
Nucleons are composed of constituent quarks and gluons; therefore, investigating the distribution of gluons, the mediators of the strong interaction, within nuclei is of fundamental importance. However, our understanding of gluon nPDFs remains markedly poorer than that of quark nPDFs. This is reflected in the significant discrepancies among gluon nPDFs extracted by different global fitting groups, especially in the intermediate-$x$ region. Whether gluons also exhibit EMC effect is an open question under active investigation Chen:2016bde ; Wang:2016mzo ; Xu:2019wso ; Hatta:2019ocp ; Wang:2020uhj ; Wang:2021elw ; Wang:2024ikx ; Yang:2023zmr ; Wang:2022kwg ; Huang:2021cac ; Wang:2024cpx ; Hu:2021fxa . Our results in the last section show that the minimum relative entropy hypothesis gives reasonable quark nPDFs. In this section, we will apply this hypothesis independently to gluon distributions.
We focus on the gluon nPDFs in the intermediate-$x$ region, which vary among different global fitting groups. Here the results from two renowned groups, EPPS21222https://research.hip.fi/qcdtheory/nuclear-pdfs/epps21/ Eskola:2021nhw and nNNPDF3.0333https://nnpdf.mi.infn.it/for-users/nnnpdf3-0/ AbdulKhalek:2022fyi are adopted. The KL divergence between the gluon nPDFs of nucleus A and deuteron is defined as
$D_{\mathrm{KL},g}^{\textrm{EMC}}(p^{A}_{g}\|q_{g}^{d})\!=\!\int_{x_{\textrm{min}}}^{x_{\textrm{max}}}\left.p_{g}^{A}(x)\right|_{\textrm{nor}}\ln\frac{\left.p_{g}^{A}(x)\right|_{\textrm{nor}}}{\left.q_{g}^{d}(x)\right|_{\textrm{nor}}}\,dx.$ (16)
Here $p_{g}^{A}(x)$ is taken as the true distribution and $q_{g}^{d}(x)$ as the reference distribution, and their values are given by global fitting groups. The limit of integration is set to $x_{\textrm{min}}=0.25$, $x_{\textrm{max}}=0.50$ for EPPS21 and $x_{\textrm{min}}=0.30$, $x_{\textrm{max}}=0.55$ for nNNPDF3.0.
These distributions necessitate normalization analogous to that in Eq. (8)
$\left\{\begin{aligned} &\left.p_{g}^{A}(x)\right|_{\textrm{nor}}=\frac{1}{\int_{x_{\textrm{min}}}^{x_{\textrm{max}}}p_{g}^{A}(x)dx}\,p_{g}^{A}(x)=z_{p,g}\,p_{g}^{A}(x)\,,\\ &\left.q_{g}^{d}(x)\right|_{\textrm{nor}}=\frac{1}{\int_{x_{\textrm{min}}}^{x_{\textrm{max}}}q_{g}^{d}(x)dx}\,q_{g}^{d}(x)=z_{q,g}\,q_{g}^{d}(x)\,.\end{aligned}\right$. (17)
The calculated values for the normalization factors $z_{p,g}$ and $z_{q,g}$ for different nuclei are collected in Appendix A.
Following the minimum relative entropy hypothesis, we determine the shape of the gluon nPDFs $\bar{p}_{g}^{A}(x)$ that minimizes the relative entropy
$D_{\mathrm{KL},g}^{\textrm{EMC}}(\bar{p}_{g}^{A}\|q_{g}^{d})=\int_{x_{\textrm{min}}}^{x_{\textrm{max}}}\!\!\left.\bar{p}_{g}^{A}(x)\right|_{\textrm{nor}}\ln\frac{\left.\bar{p}_{g}^{A}(x)\right|_{\textrm{nor}}}{\left.q_{g}^{d}(x)\right|_{\textrm{nor}}}\,dx.$ (18)
Similar to the quark case, the polynomial and canonical form of parameterizations are utilized to calculate $\bar{p}_{g}^{A}(x)$,
$\bar{p}^{A}_{g,\textrm{poly}}(x)$ $=$ $\big(a_{1}^{g}x^{3}+b_{1}^{g}x^{2}+c_{1}^{g}x+d_{1}^{g}\big)\,q_{g}^{d}(x)\,$, (19)
$\bar{p}^{A}_{g,\textrm{can}}(x)$ $=$ $\big(a_{2}^{g}x^{b_{2}^{g}}(1-x)^{c_{2}^{g}}e^{d_{2}^{g}x}\big)\,q_{g}^{d}(x)\,.$ (20)
Following the same procedure outlined in the previous section for quarks, we now directly present the results. They are shown separately according to the global fitting dataset employed.
We begin with EPPS21: the explicit results of 4He, 12C, 56Fe, 208Pb in the polynomial parametrization are
$\bar{p}^{{}^{4}\textrm{He}}_{g,\textrm{poly}}(x)$ $\!=\!$ $\big(\!-\!29.62\,x^{3}\!+\!31.38\,x^{2}\!-\!10.89\,x\!+\!2.22\big)\,q^{d}(x),$
$\bar{p}^{{}^{12}\textrm{C}}_{g,\textrm{poly}}(x)$ $\!=\!$ $\big(\!-\!36.95\,x^{3}\!+\!39.15\,x^{2}\!-\!13.58\,x\!+\!2.52\big)\,q^{d}(x),$
$\bar{p}^{{}^{56}\textrm{Fe}}_{g,\textrm{poly}}(x)$ $\!=\!$ $\big(\!-\!50.10\,x^{3}\!+\!53.05\,x^{2}\!-\!18.40\,x\!+\!3.06\big)\,q^{d}(x),$
$\bar{p}^{{}^{208}\textrm{Pb}}_{g,\textrm{poly}}(x)$ $\!=\!$ $\big(\!-\!64.57\,x^{3}\!+\!68.34\,x^{2}\!-\!23.69\,x\!+\!3.66\big)\,q^{d}(x),$
and
$\bar{p}^{{}^{4}\textrm{He}}_{\textrm{can}}(x)$ $\!=\!$ $\big(0.002\,x^{-2.56}(1-x)^{8.90}e^{21.19x}\big)\,q^{d}(x),$
$\bar{p}^{{}^{12}\textrm{C}}_{\textrm{can}}(x)$ $\!=\!$ $\big(0.0004\,x^{-3.22}(1-x)^{11.20}e^{26.68x}\big)\,q^{d}(x),$
$\bar{p}^{{}^{56}\textrm{Fe}}_{\textrm{can}}(x)$ $\!=\!$ $\big(0.003\,x^{-2.37}(1-x)^{9.83}e^{21.94x}\big)\,q^{d}(x),$
$\bar{p}^{{}^{208}\textrm{Pb}}_{\textrm{can}}(x)$ $\!=\!$ $\big(0.002\,x^{-2.36}(1-x)^{10.92}e^{23.50x}\big)\,q^{d}(x).$ (22)
for the canonical parametrization. Appendix B collects the calculated parameters ($a_{1}^{g}$, $b_{1}^{g}$, $c_{1}^{g}$ and $d_{1}^{g}$) in Eq. (V) and ($a_{2}^{g}$, $b_{2}^{g}$, $c_{2}^{g}$ and $d_{2}^{g}$) in Eq. (V) for different nuclei.
With the results in Eq. (V) and (V), we plot the gluon nPDFs given by the minimum relative entropy hypothesis in Figure 4. It can be observed that the shapes of gluon nPDFs given by the two parameterizations are highly consistent, with errors between them barely discernible. Moreover, both results agree well with the global fits of experimental data, falling within the error bar of EPPS21 result. It is worth noting that the gluon nPDF has relatively large uncertainties due to our currently poor knowledge of it. Therefore, the minimum relative entropy hypothesis may prove more useful in determining the gluon nPDFs than in constraining the quark nPDFs.
Figure 4: The obtained gluon nPDFs $\bar{p}^{A}_{g,\textrm{poly}}(x)$ with minimum KL divergence in polynomial parametrization (red line), the $\bar{p}^{A}_{g,\textrm{can}}(x)$ with minimum KL divergence in canonical parametrization (blue dashed line), as well as the gluon nPDFs $p_{g}^{A}(x)$ given by the global analysis from EPPS21 (black line).
Table 3 collects the different KL divergences. The second column lists the KL divergence calculated using the results of gluon nPDFs from EPPS21 set, while the third and fourth columns correspond to the calculated minimum KL divergence according to Eq. (V) and Eq. (V). As observed in the case of quark nPDFs, the choice of parametrization does not significantly affect the values of the minimum KL divergence.
Similarly, the $L^{2}$ norm is utilized to quantify the differences between parametrization results and the global fitting results,
$N(\bar{p}^{A}_{g,\textrm{poly}})=\sqrt{\int_{x_{\textrm{min}}}^{x_{\textrm{max}}}\Big(\bar{p}^{A}_{g,\textrm{poly}}(x)-p_{g}^{A}(x)\Big)^{2}dx}\,$, (23)
The value of norms are collected in Table 4. It can be seen that the norms given by the two parameterizations differ very little. This supports the independence of the results with respect to the forms of parameterizations. It can be observed that the $L^{2}$ norms between the gluon nPDFs derived from the minimum relative entropy hypothesis and those obtained from the EPPS21 global fits range from $1.27\times 10^{-3}$ for 3He to $1.05\times 10^{-2}$ for 208Pb, providing a quantitative measure of the discrepancy between the two.
Table 3: The KL divergences given by the global analysis from EPPS21, as well as the calculated minimum KL divergences with the polynomial and canonical forms of parameterizations in the intermediate-$x$ region $[0.25,0.50]$. The four-momentum transfer is chosen as $Q^{2}=10\,\textrm{GeV}^{2}$.
  Nucleus $D_{\mathrm{KL},g}^{\textrm{EMC}}(p^{A}_{g}\|q_{g}^{d})$ $D_{\mathrm{KL},g}^{\textrm{EMC}}(\bar{p}^{A}_{g,\textrm{poly}}\|q_{g}^{d})$ $D_{\mathrm{KL},g}^{\textrm{EMC}}(\bar{p}^{A}_{g,\textrm{can}}\|q_{g}^{d})$
3He $1.54\times 10^{-5}$ $3.34\times 10^{-6}$ $3.44\times 10^{-6}$
4He $2.31\times 10^{-4}$ $4.66\times 10^{-5}$ $4.86\times 10^{-5}$
9Be $3.24\times 10^{-4}$ $6.51\times 10^{-5}$ $7.19\times 10^{-5}$
12C $3.65\times 10^{-4}$ $33\times 10^{-5} 7.7$$70\times 10^{-5}$
56Fe $6.91\times 10^{-4}$ $1.37\times 10^{-4}$ $1.98\times 10^{-4}$
197Au $1.17\times 10^{-3}$ $2.28\times 10^{-4}$ $4.21\times 10^{-4}$
208Pb $1.19\times 10^{-3}$ $2.33\times 10^{-4}$ $3.95\times 10^{-4}$
Table 4: The norms between the gluon nPDFs derived from the minimum relative entropy hypothesis (with the second column corresponding to polynomial parameterization, the third column to canonical parameterization) and those obtained from the EPPS21 global fits.
   Nucleus   $N(\bar{p}^{A}_{g,\textrm{poly}})$   $N(\bar{p}^{A}_{g,\textrm{can}})$
3He $1.27\times 10^{-3}$ $1.25\times 10^{-3}$
4He $4.70\times 10^{-3}$ $4.58\times 10^{-3}$
9Be $5.54\times 10^{-3}$ $5.45\times 10^{-3}$
12C $5.87\times 10^{-3}$ $5.70\times 10^{-3}$
56Fe $8.02\times 10^{-3}$ $9.05\times 10^{-3}$
197Au $03\times 10^{-2} 1.1$$15\times 10^{-2}$
208Pb $1.05\times 10^{-2}$ $1.28\times 10^{-2}$
We turn next to nNNPDF3.0: applying this set of gluon global fitting data in conjunction with the minimum relative entropy hypothesis, we obtain results for 4He, 12C, 56Fe, and 208Pb within the polynomial parametrization
$\bar{p}^{{}^{4}\textrm{He}}_{g,\textrm{poly}}(x)$ $\!=\!$ $\big(\!-\!8.97\,x^{3}\!+\!10.96\,x^{2}\!-\!4.41\,x\!+\!1.60\big)\,q^{d}(x),$
$\bar{p}^{{}^{12}\textrm{C}}_{g,\textrm{poly}}(x)$ $\!=\!$ $\big(\!-\!47.08\,x^{3}\!+\!57.52\,x^{2}\!-\!23.11\,x\!+\!4.13\big)\,q^{d}(x),$
$\bar{p}^{{}^{56}\textrm{Fe}}_{g,\textrm{poly}}(x)$ $\!=\!$ $\big(\!-\!216.67\,x^{3}\!+\!263.53\,x^{2}\!-\!105.39\,x\!+\!15.07\big)\,q^{d}(x),$
$\bar{p}^{{}^{208}\textrm{Pb}}_{g,\textrm{poly}}(x)$ $\!=\!$ $\big(\!-\!389.23\,x^{3}\!+\!470.88\,x^{2}\!-\!187.19\,x\!+\!25.70\big)\,q^{d}(x),$
and
$\bar{p}^{{}^{4}\textrm{He}}_{\textrm{can}}(x)$ $\!=\!$ $\big(0.0976\,x^{-1.03}(1-x)^{2.22}e^{6.33x}\big)\,q^{d}(x),$
$\bar{p}^{{}^{12}\textrm{C}}_{\textrm{can}}(x)$ $\!=\!$ $\big(0.0014\,x^{-2.86}(1-x)^{7.16}e^{19.15x}\big)\,q^{d}(x),$
$\bar{p}^{{}^{56}\textrm{Fe}}_{\textrm{can}}(x)$ $\!=\!$ $\big(0.0025\,x^{-2.20}(1-x)^{15.47}e^{30.52x}\big)\,q^{d}(x),$
$\bar{p}^{{}^{208}\textrm{Pb}}_{\textrm{can}}(x)$ $\!=\!$ $\big(0.0002\,x^{-2.60}(1-x)^{31.26}e^{56.80x}\big)\,q^{d}(x).$
for the canonical parametrization. Appendix B collects the calculated parameters ($a_{1}^{g}$, $b_{1}^{g}$, $c_{1}^{g}$ and $d_{1}^{g}$) in Eq. (V) and ($a_{2}^{g}$, $b_{2}^{g}$, $c_{2}^{g}$ and $d_{2}^{g}$) in Eq. (V).
The gluon nPDFs obtained from the minimum relative entropy hypothesis are shown in Fig. 5. Interestingly, these nPDFs begin to exhibit noticeable deviations from the nNNPDF3.0 results, with the discrepancy becoming particularly pronounced for 56Fe and 208Pb. Their corresponding KL divergences are shown in Table 5. This discrepancy is more clearly illustrated by computing the $L^{2}$ norms between the gluon nPDFs derived from the minimum relative entropy hypothesis and those obtained from the nNNPDF3.0 global fits. These values are presented in Table 6, where the norms are generally larger than those obtained with EPPS21 previously. For example, the results for 56Fe, 197Au, and 208Pb are approximately four times larger than their EPPS21 counterparts. If the minimum relative entropy hypothesis is taken as a benchmark for evaluating global fits of gluon nPDFs, the EPPS21 results align more favorably with this criterion. This provides a novel information-theoretic perspective for investigating gluon nPDFs.
Figure 5: The obtained gluon nPDFs $\bar{p}^{A}_{g,\textrm{poly}}(x)$ with minimum KL divergence in polynomial parametrization (red line), the $\bar{p}^{A}_{g,\textrm{can}}(x)$ with minimum KL divergence in canonical parametrization (blue dashed line), as well as the gluon nPDFs $p_{g}^{A}(x)$ given by the global analysis from nNNPDF3.0 (black line).
Table 5: The KL divergences given by the global analysis from nNNPDF3.0, as well as the calculated minimum KL divergences with the polynomial and canonical form of parameterizations in the intermediate-$x$ region $[0.30,0.55]$. The four-momentum transfer is chosen as $Q^{2}=10\,\textrm{GeV}^{2}$.
  Nucleus $D_{\mathrm{KL},g}^{\textrm{EMC}}(p^{A}_{g}\|q_{g}^{d})$ $D_{\mathrm{KL},g}^{\textrm{EMC}}(\bar{p}^{A}_{g,\textrm{poly}}\|q_{g}^{d})$ $D_{\mathrm{KL},g}^{\textrm{EMC}}(\bar{p}^{A}_{g,\textrm{can}}\|q_{g}^{d})$
4He $4.55\times 10^{-5}$ $4.27\times 10^{-6}$ $4.30\times 10^{-6}$
9Be $5.23\times 10^{-4}$ $5.34\times 10^{-5}$ $6.93\times 10^{-5}$
12C $1.02\times 10^{-3}$ $1.07\times 10^{-4}$ $1.50\times 10^{-4}$
56Fe $1.45\times 10^{-2}$ $1.93\times 10^{-3}$ $5.30\times 10^{-3}$
197Au $3.69\times 10^{-2}$ $5.49\times 10^{-3}$ $1.66\times 10^{-2}$
208Pb $4.38\times 10^{-2}$ $6.53\times 10^{-3}$ $2.21\times 10^{-2}$
Table 6: The norms between the gluon nPDFs derived from the minimum relative entropy hypothesis (with the second column corresponding to polynomial parameterization, the third column to canonical parameterization) and those obtained from the nNNPDF3.0 global fits.
   Nucleus   $N(\bar{p}^{A}_{g,\textrm{poly}})$   $N(\bar{p}^{A}_{g,\textrm{can}})$
4He $1.69\times 10^{-3}$ $1.85\times 10^{-3}$
9Be $5.97\times 10^{-3}$ $6.12\times 10^{-3}$
[…]
208Pb $4.43\times 10^{-2}$ $6.59\times 10^{-2}$
The results in this work are intriguing which worth further investigation to uncover the profound physical connections between PDFs and nPDFs within the framework of quantum information theory. The behavior of nonperturbative QCD is very complex. In contrast to conventional approaches, the idea of entropy offers a promising avenue for elucidating the fundamental symmetries and underlying dynamics of QCD that are otherwise hidden. The present work also has aspects that require further improvement. For instance, extending the applicable range from the EMC region to the entire region remains unknown — we leave such questions to future study.
VI Concluding remarks
In this work, beyond conventional observables, we have introduced the KL divergence, which is a well-known estimators of quantum information theory, as a measure to probe the parton structure of nucleon bound in nucleus. The obtained results of quark nPDFs are consistent with the latest global fits of experimental data, and the proposed minimum relative entropy hypothesis can serve as a benchmark for assessing the quality of global fits to gluon nPDFs. These findings have revealed the promising potential for the interdisciplinary research to parton distributions.
In particle physics, determination of nonperturbative inputs apart from the global fitting is not only a complementarity but also a necessity to understand the nature of hadrons. The introductions of KL divergence and minimum relative entropy hypothesis offer novel perspective on old problems. It can be used for searching more details of the nuclear structure function and parton distributions. Admittedly, all the results presented in this work should rather be considered primarily as indicative evidence; it is hoped that this work could inspire further research on applying information theory to nucleon structure studies.
Acknowledgements
The authors would like to thank Prof. Wei Wang and Shuai Zhao for fruitful and inspiring discussions. This work is supported in part by the National Natural Science Foundation of China under Grant No. 12335003, and by the Fundamental Research Funds for the Central Universities under No. lzujbky-2023-stlt01, lzujbky-2024-oy02 and lzujbky-2025-eyt01, the Scientific Research Innovation Capability Support Project for Young Faculty under Grant No. ZYGXQNJSKYCXNLZCXM-P2. J.X is supported in part by the National Natural Science Foundation of China under Grant No. 12475098 and 12105247.
Appendix A Different $x$-region KL divergence calculation details
Here, we compute the KL divergence from the Eqs. (8) and (9). The corresponding integration limits are $10^{-6},0.99]$ and $[0.25,0.65]$. The resulting KL divergences are designated as $D_{\mathrm{KL}}^{\textrm{full}}(p^{A}\|q^{d})$ and $D_{\mathrm{KL}}^{\textrm{EMC}}(p^{A}\|q^{d})$. Table 7 presents the calculated KL divergences for different nuclei.
Table 7: The KL divergences for different nuclei in full-$x$ region $10^{-6},0.99]$ and intermediate-$x$ region $[0.25,0.65]$. The four-momentum transfer is chosen as $Q^{2}=10\,\textrm{GeV}^{2}$.
   Nucleus   $D_{\mathrm{KL}}^{\textrm{full}}(p^{A}\|q^{d})$   $D_{\mathrm{KL}}^{\textrm{EMC}}(p^{A}\|q^{d})$
3He $1.33\times 10^{-3}$ $0.46\times 10^{-4}$
4He $5.00\times 10^{-3}$ $2.32\times 10^{-4}$
9Be $5.22\times 10^{-3}$ $4.75\times 10^{-4}$
12C $5.42\times 10^{-3}$ $3.70\times 10^{-4}$
56Fe $6.38\times 10^{-3}$ $8.17\times 10^{-4}$
197Au $7.93\times 10^{-3}$ $16.19\times 10^{-4}$
208Pb $8.03\times 10^{-3}$ $16.85\times 10^{-4}$
Here, we utilize the nPDF sets from EPPS21 and CT18A. For full-$x$ region $10^{-6},0.99]$ and intermediate-$x$ region $[0.25,0.65]$, we take the $F_{2}^{A}(x,Q^{2})/A$ as $p^{A}(x)$ in Eq. (6) and the $F_{2}^{d}(x,Q^{2})/2$ as $q^{d}(x)$. Qualitatively speaking, as A increases, nuclear modification becomes more apparent, leading to an increase in the KL divergence—a trend that is consistent with the results shown in Table 7. However, it is worth noting that the value for $^{12}\textrm{C}$} is lower than for $^{9}\textrm{Be}$}, a pattern consistent with findings in studies of nucleon short-range correlations which show a non-monotonic dependence on A for $^{12}\textrm{C}$} and $^{9}\textrm{Be}$} CLAS:2019vsb . This may arise from the unique nature of beryllium, whose density is similar to helium, yet whose EMC effect closely resembles that of denser nuclei due to its significant cluster structure.
The values of normalization factors $z_{p}$ and $z_{q}$ defined in Eq. (8) are listed in Table 8.
Table 8: The normalization factors $z_{p}$ and $z_{q}$ for different nuclei in full-$x$ region $10^{-6},0.99]$ and intermediate-$x$ region $[0.25,0.65]$. The four-momentum transfer is chosen as $Q^{2}=10\,\textrm{GeV}^{2}$.
   Nucleus   $z_{p}^{\textrm{full}}$   $z_{p}^{\textrm{EMC}}$
2D($z_{q}$) $7.06$ $19.61$
3He $6.70$ $18.16$
4He $7.12$ $20.32$
9Be $7.26$ $21.05$
12C $7.14$ $20.48$
56Fe $7.27$ $21.16$
197Au $7.48$ $22.20$
208Pb $7.50$ $22.29$
Table 9 presents the normalization factors used in the calculation of the gluon KL divergence.
Table 9: The normalization factors $z_{p,g}$ and $z_{q,g}$ for different nuclei in intermediate-$x$ region $[0.25,0.5]$. The four-momentum transfer is chosen as $Q^{2}=10\,\textrm{GeV}^{2}$.
   Nucleus   $z_{p,g}^{\textrm{EMC}}$
2D($z_{q,g}$) $7.67$
3He $7.74$
4He $7.93$
9Be $7.97$
12C $7.98$
56Fe $8.09$
197Au $8.20$
208Pb $8.21$
Table 10 presents the normalization factors used in the calculation of the nNNPDF30 gluon KL divergence .
Table 10: The normalization factors $z_{p,g}$ and $z_{q,g}$ for different nuclei in intermediate-$x$ region $[0.3,0.55]$. The four-momentum transfer is chosen as $Q^{2}=10\,\textrm{GeV}^{2}$.
   Nucleus   $z_{p,g}^{\textrm{EMC}}$
2D($z_{q,g}$) $13.03$
4He $12.99$
9Be $12.79$
12C $12.65$
56Fe $11.94$
197Au $12.32$
208Pb $12.48$
Appendix B Collection of determined parameters
In this appendix, we collect the determined parameters in the polynomial and canonical parameterizations. For the quark case, we give the four parameters ($a_{1}$, $b_{1}$, $c_{1}$ and $d_{1}$) in Eq. (11) and four parameters ($a_{2}$, $b_{2}$, $c_{2}$ and $d_{2}$) in Eq. (13), which are determined by endpoints constrains and minimizing the KL divergences in Eq. (10) for different nuclei.
Table 11: For the quark case, the determined parameters in polynomial form in Eq. (11) and canonical form in Eq. (13) for different nuclei. The integration interval is $x\in[0.25,0.65]$.
   Nucleus     $a_{1}$     $b_{1}$     $c_{1}$     $d_{1}$     $a_{2}$     $b_{2}$     $c_{2}$     $d_{2}$
3He $4.29$ 5.39-5.39 $2.19$ $0.79$ $2.99$ $0.45$ 0.90-0.90 2.68-2.68
4He 5.89-5.89 $7.39$ 3.00-3.00 $1.38$ $0.20$ 0.69-0.69 $1.39$ $4.13$
9Be 8.75-8.75 $10.97$ 4.45-4.45 $1.54$ $0.08$ 1.06-1.06 $2.15$ $6.37$
12C 7.60-7.60 $9.53$ 3.86-3.86 $1.48$ $0.13$ 0.90-0.90 $1.81$ $5.37$
56Fe 11.51-11.51 $14.43$ 5.85-5.85 $1.72$ $0.039$ 1.40-1.40 $2.84$ $8.43$
197Au 15.94-15.94 $19.97$ 8.08-8.08 $1.98$ $0.009$ 2.04-2.04 $4.16$ $12.30$
208Pb 16.23-16.23 $20.32$ 8.22-8.22 $1.99$ $0.008$ 2.09-2.09 $4.25$ $12.57$
For the gluon case, after applying the EPPS21 global fitting for gluon nPDFs and the minimum relative entropy hypothesis, we give the four parameters ($a_{1}^{g}$, $b_{1}^{g}$, $c_{1}^{g}$ and $d_{1}^{g}$) in Eq. (V) and four parameters ($a_{2}^{g}$, $b_{2}^{g}$, $c_{2}^{g}$ and $d_{2}^{g}$) in Eq. (V), which are determined by endpoints constrains and minimizing the KL divergences in Eq. (18) for different nuclei.
Table 12: For the gluon case, the determined parameters in polynomial form in Eq. (19) and canonical form in Eq. (20) for different nuclei by utilizing the EPPS21 global fits. The integration interval is $x\in[0.25,0.50]$.
   Nucleus     $a_{1}^{g}$     $b_{1}^{g}$     $c_{1}^{g}$     $d_{1}^{g}$     $a_{2}^{g}$     $b_{2}^{g}$     $c_{2}^{g}$     $d_{2}^{g}$
3He 8.09-8.09 $8.57$ 2.98-2.98 $1.33$ $0.20$ 0.65-0.65 $2.29$ $5.43$
4He 29.62-29.62 $31.38$ 10.89-10.89 $2.22$ $0.002$ 2.56-2.56 $8.90$ $21.19$
9Be 34.85-34.85 $36.92$ 12.81-12.81 $2.44$ $0.002$ 2.49-2.49 $9.06$ $21.22$
12C 36.95-36.95 $39.15$ 13.58-13.58 $2.52$ $0.0004$ 3.22-3.22 $11.20$ $26.68$
56Fe 50.10-50.10 $53.05$ 18.40-18.40 $3.06$ $0.003$ 2.37-2.37 $9.83$ $21.94$
197Au 60.90-60.90 $67.63$ 23.44-23.44 $3.63$ $0.006$ 2.00-2.00 $9.40$ $20.05$
208Pb 64.57-64.57 $68.34$ 23.69-23.69 $3.66$ $0.002$ 2.36-2.36 $10.92$ $23.50$
For the gluon case, after applying the nNNPDF3.0 global fitting for gluon nPDFs and the minimum relative entropy hypothesis, we give the four parameters ($a_{1}^{g}$, $b_{1}^{g}$, $c_{1}^{g}$ and $d_{1}^{g}$) in Eq. (V) and four parameters ($a_{2}^{g}$, $b_{2}^{g}$, $c_{2}^{g}$ and $d_{2}^{g}$) in Eq. (V), which are determined by endpoints constrains and minimizing the KL divergences in Eq. (18) for different nuclei.
Table 13: For the gluon case, the determined parameters in polynomial form in Eq. (V) and canonical form in Eq. (V) for different nuclei by utilizing the nNNPDF3.0 global fits. The integration interval is $x\in[0.3,0.55]$.
   Nucleus     $a_{1}^{g}$     $b_{1}^{g}$     $c_{1}^{g}$     $d_{1}^{g}$     $a_{2}^{g}$     $b_{2}^{g}$     $c_{2}^{g}$     $d_{2}^{g}$
4He 8.97-8.97 $10.96$ 4.41-4.41 $1.60$ $0.0976$ 1.03-1.03 $2.22$ $6.33$
9Be 32.70-32.70 $39.96$ 16.06-16.06 $3.17$ $0.0060$ 2.24-2.24 $5.44$ $14.74$
12C 47.08-47.08 $57.52$ 23.11-23.11 $4.13$ $0.0014$ 2.86-2.86 $7.16$ $19.15$
56Fe 216.67-216.67 $263.53$ 105.39-105.39 $15.07$ $0.0025$ 2.20-2.20 $15.47$ $30.52$
197Au 359.78-359.78 $435.71$ 173.42-173.42 $23.93$ $1.58*10^{-6}$ 4.96-4.96 $30.91$ $62.49$
208Pb 389.23-389.23 $470.88$ 187.19-187.19 $25.70$ $0.0002$ 2.60-2.60 $31.26$ $56.80$